I need to know the radius and center for this 
equation. It is for circles: 

256x² + 256y² = 64


Answer by AnlytcPhil(1806)

I need to know the radius and center for this 
equation. It is for circles. 
256x² + 256y² = 64.

The equation of a circle is

(x - h)² + (y - k)² = r²

where the center is (h,k) and the radius is r.

      256x² + 256y² = 64

Divide thru by 256

    256x²     256y²     64
   ------- + ------ = -----
     256       256     256

     1         1         1
    25625664
   ------- + ------ = -----
     256       256     256
      1         1       4

            x² + y² = 1/4

Write x as (x - 0) and y as (y - 0)

(x - 0)² + (y - 0)² = 1/4

Compare to

(x - h)² + (y - k)² = r²

so h = 0, k = 0, r² = 1/4, so r = 1/2

Center = (h, k) = (0, 0), radius = 1/2

Edwin
AnlytcPhil@aol.com